# On the equivalence of all notions of generalized derivations whose domain is a C-algebra

### v1

@@ 3. Automatic continuity of generalized Jordan derivation of the third type
Assume next that $\mathcal{M}=\mathcal{A}^{*}$. A consequence of the Grothendieck’s inequality established by Haagerup in [16] asserts that every bounded linear mapping from a C-algebra to the dual space of another C-algebra factors through a Hilbert space, and thus it is weakly compact. Therefore, by Gantmacher’s theorem [14], $F^{**}\left(\mathcal{A}^{**}\right)\subseteq\mathcal{M}^{*}$ and $G^{**}\left(\mathcal{A}^{**}\right)\subseteq\mathcal{M}^{*}$. We claim that
@@
We can conclude now that for a non-commutative C-algebra $\mathcal{A}$, the class of generalized Jordan derivations of the third type on $\mathcal{A}$ is strictly included in the class of generalized Jordan derivation of the second type. Consider, for example, an element $\zeta$ which is not in the centre of $\mathcal{A}$ and the mapping $F:\mathcal{A}\to\mathcal{A}$, $F(a)=a\circ\zeta$ which is a generalized Jordan derivation of the second type but not of the third type.
Finally, it is worth to say a few words about what could be the Jordan version of generalized derivations of the first type. Let $\mathcal{M}$ be a Banach $\mathcal{A}$-bimodule on a Banach algebra $\mathcal{A}$. A generalized Jordan derivation of the first type is a linear mapping $F:\mathcal{A}\to\mathcal{M},$ for which there exists a Jordan derivation $D:\mathcal{A}\to\mathcal{M}$ satisfying $F(a\circ b)=F(a)\circ b+a\circ D(b)$ for all $a,b\in\mathcal{A}$. Indeed some algebraists defined the so-called “Jordan generalized derivation" by letting $D$ to be just a linear map. That is the case in the paper by Li and Benkovič [33], where it is shown that any generalized Jordan derivation of the first type on a triangular algebra is a kind of generalized derivation of the third type. Every Jordan derivation of the first type is automatically a generalized Jordan derivation of the third type, and thus continuous by Theorem 3.1 when $\mathcal{A}$ is a C-algebra. Furthermore, if $\mathcal{M}$ is essential, the sets of all generalized Jordan derivations of the first and third type, respectively, are strictly included in the set of all generalized Jordan derivation of the second type, and actually they are all generalized derivations of the first/second/third type.
Despite we have tried to clarify the relationships between the different notions of “generalized derivations”, several interesting questions remain open after our study. Can we remove the hypothesis $\mathcal{M}$ being essential in Proposition 3.2? Is the final conclusion in the same proposition true for other Banach bimodules appart from $\mathcal{A}$ and $\mathcal{A}^{*}$?
It is also natural to expect that the notions of generalized derivations of the third type studied in this note give rise to appropriate concepts in the setting of JB-algebras and JB-triples. In the case of a unital JB-algebra $\mathcal{J}$,the techniques and arguments in [17] can be appropriately modified to prove that every generalized Jordan derivation of the third type on $\mathcal{J}$ or from $\mathcal{J}$ to $\mathcal{J}^{*}$ is automatically continuous. However, under more general hypotheses the answer is not clear.

### v2

@@ 3. Automatic continuity of generalized Jordan derivation of the third type
Assume next that $\mathcal{M}=\mathcal{A}^{*}$. A consequence of the Grothendieck’s inequality, established by Haagerup in [17, page 95], asserts that every bounded linear mapping from a C-algebra to the dual space of another C-algebra factors through a Hilbert space, and thus it is weakly compact. Therefore, by Gantmacher’s theorem [15], $F^{**}\left(\mathcal{A}^{**}\right)\subseteq\mathcal{M}^{*}$ and $G^{**}\left(\mathcal{A}^{**}\right)\subseteq\mathcal{M}^{*}$. We claim that
@@
We can conclude now that for a non-commutative C-algebra $\mathcal{A}$, the class of generalized Jordan derivations of the third type on $\mathcal{A}$ is strictly included in the class of generalized Jordan derivations of the second type. Consider, for example, an element $\zeta$ which is not in the centre of $\mathcal{A}$ and the mapping $F:\mathcal{A}\to\mathcal{A}$, $F(a)=a\circ\zeta$ which is a generalized Jordan derivation of the second type but not of the third type.
Finally, it is worth to say a few words about what could be the Jordan version of generalized derivations of the first type. Let $\mathcal{M}$ be a Banach $\mathcal{A}$-bimodule on a Banach algebra $\mathcal{A}$. A generalized Jordan derivation of the first type is a linear mapping $F:\mathcal{A}\to\mathcal{M},$ for which there exists a Jordan derivation $D:\mathcal{A}\to\mathcal{M}$ satisfying $F(a\circ b)=F(a)\circ b+a\circ D(b)$ for all $a,b\in\mathcal{A}$. Indeed some algebraists defined the so-called “Jordan generalized derivation" by letting $D$ be just a linear map. That is the case in the paper by Li and Benkovič [34], where it is shown that any generalized Jordan derivation of the first type on a triangular algebra is a kind of generalized derivation of the third type. Every Jordan derivation of the first type is automatically a generalized Jordan derivation of the third type, and thus continuous by Theorem 3.1 when $\mathcal{A}$ is a C-algebra.
Despite we have tried to clarify the relationships between the different notions of “generalized derivations”, several interesting questions remain open after our study. Can we remove the hypothesis $\mathcal{M}$ being essential in Proposition 3.2? Is the final conclusion in the same proposition true for other Banach bimodules besides $\mathcal{A}$ and $\mathcal{A}^{*}$?
It is also natural to expect that the notions of generalized derivations of the third type studied in this note give rise to appropriate concepts in the setting of JB-algebras and JB-triples. In the case of a unital JB-algebra $\mathcal{J}$,the techniques and arguments in [18] can be appropriately modified to prove that every generalized Jordan derivation of the third type on $\mathcal{J}$ or from $\mathcal{J}$ to $\mathcal{J}^{*}$ is automatically continuous. However, under weaker hypotheses the conclusion is not so clear.
@@
Disclosure: All authors declare that they have no conflicts of interest to disclose.